OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Ocean Tours. Age Ticket Price Under three years free Three to twelve years $36 Over twelve years $48 On a certain day, the company took 92 people on whale watching trips. There were 40 children aged twelve and under, of which some children were under three years. If x represents the number of children undr three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day? (92 - x)36 + (92 – 40) ∙ 48 = C (40 + x)36 + (92 – (40 + x))

OpenStudy (anonymous):

(92 - x)36 + (92 – 40) ∙ 48 = C (40 + x)36 + (92 – (40 + x)) ∙ 48 = C (92 + x)36 + (92 – (40 + x)) ∙ 48 = C (40 – x)36 + (92 – 40) ∙ 48 = C

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

It's the last selection because 40 - x children are chargeable and there are 92 - 40 adults paying top admission.

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @Para_phrenic88